OpenStudy (anonymous):

Kevin is trying to find a white sock in his drawer. He has 16 white socks, 4 brown socks, and 6 black socks. What is the probability that he pull out either a black or brown sock, puts it back, and then pulls out a white sock? A)9/13 B)20/13 C)40/169 D)96/169

OpenStudy (yanasidlinskiy):

there are 26 socks in total (16 + 4 + 6) so we are asked if if he pulls out a black OR white sock. the or should signal to you that you add these probabilities \(16/26 + 6/26 = 22/26\) now, it says that he puts it back in (in math terms, he replaces it) and then picks out a brown sock \(4/26\) now we just multiply those two probabilities together. the key word you want to look for is AND. and implies that you multiply the probabilites \(22/26*4/26 = 22/13*1/13 = 22/169\)

OpenStudy (anonymous):

I thought that was it! Thank you so much cx

OpenStudy (yanasidlinskiy):

Welcome!!:)